# Coarse graining correlation matrices according to macrostructures: Financial markets as a paradigm

### v1

@@ A.1 S&P 500
Figure A.4: Transition matrices showing transitions between different market states for S&P 500 in Fig. 6 obtained with (a) Pearson correlation matrices and (b) sectorial Guhr matrices. The transition matrices are near-tridiagonal and the necessary criterion for Markovianity given in Eq. (2) of [6] is fulfilled, for both. The equilibrium distributions corresponding to (a) and (b) are (0.307, 0.273, 0.236, 0.123, 0.061) and (b) (0.272, 0.297, 0.229, 0.147, 0.055), respectively.
@@ A.2 Nikkei 225
Figure A.8: Transition matrices showing transitions between different market states for Nikkei 225 in Fig. A.4 obtained with (a) Pearson correlation matrices and (b) sectorial Guhr matrices. The transition matrices are near-tridiagonal and the necessary criterion for Markovianity given in Eq. (2) of [6] is fulfilled, for both. The equilibrium distributions corresponding to (a) and (b) are $(0.198,0.168,0.257,0.087,0.189,0.100)$ and $(0.155,0.270,0.113,0.206,0.163,0.093)$ respectively.

### v2

@@ A.1 S&P 500
Figure A.4: Transition matrices showing transitions between different market states for S&P 500 in Fig. 6 obtained with (a) Pearson correlation matrices and (b) sectorial Guhr matrices. The transition matrices are near-tridiagonal and the necessary criterion for Markovianity given in Eq. (2) of Pharasi et al. (2018) is fulfilled, for both. The equilibrium distributions corresponding to (a) and (b) are (0.307, 0.273, 0.236, 0.123, 0.061) and (b) (0.272, 0.297, 0.229, 0.147, 0.055), respectively.
@@ A.2 Nikkei 225
Figure A.8: Transition matrices showing transitions between different market states for Nikkei 225 in Fig. A.4 obtained with (a) Pearson correlation matrices and (b) sectorial Guhr matrices. The transition matrices are near-tridiagonal and the necessary criterion for Markovianity given in Eq. (2) of Pharasi et al. (2018) is fulfilled, for both. The equilibrium distributions corresponding to (a) and (b) are $(0.198,0.168,0.257,0.087,0.189,0.100)$ and $(0.155,0.270,0.113,0.206,0.163,0.093)$ respectively.